### v2

@@ Abstract
Context. Gamma-ray bursts (GRBs), observed at redshifts as high as 9.4, could serve as valuable probes for investigating the distant Universe. However, this necessitates an increase in the number of GRBs with determined redshifts, as currently, only 12% of GRBs have known redshifts due to observational biases.
Aims. We aim to address the shortage of GRBs with measured redshifts, enabling us to fully realize their potential as valuable cosmological probes
Methods. Following Dainotti et al. (2024c), we have taken a second step to overcome this issue by adding 30 more GRBs to our ensemble supervised machine learning training sample, an increase of 20%, which will help us obtain more accurate pseudo-redshifts. In addition, we have built a freely accessible and user-friendly web app that infers the redshift of long GRBs (LGRBs) with plateau emission using our machine learning model. The web app is the first of its kind for such a study and will allow the community to obtain pseudo-redshifts by entering the GRB parameters in the app.
Results. Through our machine learning model, we have successfully estimated redshifts for 276 LGRBs using X-ray afterglow parameters detected by the Neil Gehrels Swift Observatory and increased the sample of LGRBs with known redshifts by 110%. We also perform Monte Carlo simulations to demonstrate the future applicability of this research.
Conclusions. The results presented in this research will enable the community to increase the sample of GRBs with known pseudo-redshifts. This can help address many outstanding issues, such as GRB formation rate, luminosity function, and the true nature of low-luminosity GRBs, and enable the application of GRBs as standard candles.
@@ 1 Introduction
Gamma-Ray Bursts (GRBs) are the brightest and most powerful explosion events after the Big Bang and are observed across a wide range of redshifts, from 0.0085 (Galama et al., 1998) to 8.2 and 9.4 (Tanvir et al., 2008; Cucchiara et al., 2011). This vast range of redshifts ($z$) makes GRBs ideal probes for studying the early Universe and tracking its evolution, providing insights into the composition, expansion rate, and other key aspects. GRBs, generally detected in $\gamma$-rays, X-rays, and occasionally in optical wavelengths, are classified into two types based on their duration, characterized by the $T_{90}$ parameter. $T_{90}$ is the time span during which the GRB emits 90% of its total observed fluence in $\gamma$-rays. GRBs with $T_{90}<2s$ or $T_{90}>2s$ and are classified as short GRBs (SGRBs) and long GRBs (LGRBs), respectively.
LGRBs are produced from core collapse supernovae (Mazets et al., 1981; Kouveliotou et al., 1993; Paczyński, 1998; Woosley & Bloom, 2006), while SGRBs are produced by neutron star-neutron star or neutron star-black hole mergers (Narayan et al., 1992; Duncan & Thompson, 1992; Usov, 1992; Thompson, 1994; Nakar, 2007; Goldstein et al., 2017a, b; Abbott et al., 2017).
@@
In addition to these two main categories of GRBs, numerous other sub-classes have been identified in the literature. These classes are based only on observational properties such as the $T_{90}$ duration, spectral, and lightcurve properties. These include classes such as Short GRBs with extended emission (SEE, Norris et al. (2000); Norris & Bonnell (2006); Levan et al. (2007); Norris et al. (2010)), where $T_{90}>2$ seconds but with similar spectral properties to SGRBs, Very Long GRBs (VLGRBs, Levan et al. (2014)) and Ultra Long GRBs (ULGRBs, Stratta et al. (2013); Zhang et al. (2014); Gendre et al. (2019) where $T_{90}>500$ and $1000$ seconds, respectively. Zhang et al. (2007) unified these various classes based on their GRB progenitors. According to their prescription, Type-I GRBs comprise of classes which are SGRBs, SEE, and others similar to SGRBs, and occur due to compact binary mergers (Nakar, 2007; Abbott et al., 2017). Type-II GRBs comprise LGRBs, VLGRBs, ULGRBs, and other similar classes with core-collapse supernova progenitors (the so-called collapsar events) (Woosley, 1993; Woosley et al., 1993). Irrespective of these efforts, an exhaustive classification of GRBs has not yet been found, and this is an active field of research.
@@
Phenomenologically, the GRB lightcurve (LC) from across all the classes can be separated into two distinct phases. The prompt emission is the main emission, characterized by a burst of high-energy photons spanning from $\gamma$-rays to hard X-rays and sometimes extending into optical wavelengths (Vestrand et al., 2005; Beskin et al., 2010; Gorbovskoy et al., 2012; Vestrand et al., 2014). Following the prompt, the afterglow is a long-lasting, multiwavelength emission powered by the interaction of the GRB ejecta with the surrounding medium (van Paradijs et al., 1997; Costa et al., 1997; Piro et al., 1998). It is detected across various wavelengths, such as X-rays, optical, and occasionally even radio.
@@
A peculiar trait of GRBs, uncovered by the Neil Gehrels Swift observatory (hereafter Swift, Gehrels et al. (2004)), is the presence of a flat region in the LC, the so-called plateau emission, which is characterized by almost constant flux values (Nousek et al., 2006; Zhang et al., 2006; O’Brien et al., 2006; Sakamoto et al., 2007; Liang et al., 2007b; Dainotti et al., 2008; Zaninoni et al., 2013; Rowlinson et al., 2014a). The end of the plateau is considered to be a point in the LC’s afterglow, where the luminosity begins to decay with time as a powerlaw. The plateau feature has been found in 42% of X-ray afterglows (Evans et al., 2009; Li et al., 2018). In our recent estimate, the percentage of GRBs with observable plateaus is 38.5%.
Currently, the primary hurdle in the LGRB population study is the lack of a large GRB sample with known redshift. Quick localization and spectral data are required to determine GRBs’ redshift.
@@
Swift allows for this quick detection and multi-wavelength follow-up in $\gamma$-rays, X-rays, and ultraviolet/optical. It has three telescopes: the Burst Alert Telescope (BAT, Burrows et al. (2005)), the X-ray Telescope (XRT, Barthelmy et al. (2005)), and the Ultraviolet/Optical Telescope (UVOT, Roming et al. (2005)). Swift uses XRT for precise localization, and with UVOT, it obtains spectra and sometimes redshifts. Swift’s rapid localization abilities have significantly contributed to exploring the high-$z$ Universe.
@@
However, although Swift triggers many observation programs, only 26% (431) of Swift’s 1656 GRBs have known spectroscopic redshifts (as of 20th August 2024). The Fermi Telescope has observed more than 3700 GRBs, with its Large Area Telescope (LAT, Ajello et al. (2019)) and the Fermi-Gamma-Ray Burst Monitor (GBM, Meegan et al. (2009)). Currently, only 12% of GRBs detected by Swift and Fermi have spectroscopic redshift ($z_{obs}$), $\sim$90of which are observed by Swift.
Measuring redshift, especially high-$z$, is challenging due to limited telescope time and the difficulty of securing time for follow-up observations using large aperture telescopes. Therefore, determining the $z$ of GRBs via alternative methods is crucial.
@@
Correctly determining the GRB luminosity function (LF) is one of the main goals that can be achieved by increasing the number of GRBs with known redshifts. The LF gives us the number of GRBs per unit luminosity and is vital for understanding the properties of the GRB population, such as their energy release and emission mechanisms. Another principal objective is determining the cosmic GRB formation rate (GRBFR). The GRBFR measures the frequency of GRBs across cosmic time.
@@
Petrosian et al. (2015); Petrosian & Dainotti (2024) have found a discrepancy at low redshift ($10<z<$ 1) between the LGRB formation rate and the star formation rate (SFR). This discrepancy, which has been observed by several research groups independently, remains an active area of investigation. Addressing this requires a bigger sample of GRBs with known redshift, thus highlighting the importance of our study.
@@
The correct determination of the GRB rate for both LGRBs and SGRBs is hampered by the broad range of GRB physical features; therefore, a precise distribution of these features and their evolution over cosmic time is necessary. Some of us investigated this issue with a sample of 179 GRBs with optical LCs (Dainotti et al., 2024b), where we determined the luminosity and formation rate evolutions and the general shape of the LF after correcting for evolution. We observed the average LGRB’s rate at low-$z$ ($z<1$) to be 10.46 Gpc-3 $yr^{-1}$, which is roughly 3 times more than Hopkins & Beacom (2006) SFR (3.7 Gpc-3 $yr^{-1}$). This LGRB rate is an actual rate where the normalization is in physical units, which contrasts with the arbitrary normalization used in most current literature.
@@
To address the challenges outlined above, we propose here a machine learning (ML) based approach for estimating the redshift of GRBs. This work is a follow-up of Dainotti et al. (2024c) and Dainotti et al. (2024b) and utilizes an enlarged sample, 20% larger than the sample in Dainotti et al. (2024c), to train our ML models. In addition, we have extended the generalization sample by 35%. This is the set of GRBs without known redshifts, and we predict the redshift of 276 GRBs using our trained ML model. Furthermore, we release a user-friendly web application that provides pseudo-redshifts for new GRBs, built on the ML models described in this paper.
@@
The paper is structured as follows: Sec. 2 discusses the advantages of using an ML approach compared to using linear relations for the redshift estimation. In Sec. 3, we describe our dataset. In Sec. 4, we describe the functionalities and various modules of our web app, as well as details of the ML models being used. We also describe the Monte Carlo (MC) simulation performed to understand the viability of our ML models in relation to future data. Sec. 5 shows the results obtained from our analysis and the usability of the web app. Here, we also present the future applicability of our ML approach. In Sec. 6, we discuss the relevance of our results. Finally, we summarize and conclude in Sec. 7.
2 Advantage of machine learning VS linear relationships
@@
Expanding the GRB sample with redshift provides a significant benefit, notably the possibility of using them as standardized candles, objects with known intrinsic luminosities, or with luminosities that can be derived via established relationships among GRB properties. The adoption of empirical relations between the distance-dependent and intrinsic properties of GRBs enables cosmological studies in unexplored redshift ranges. One of the earliest efforts is the Dainotti Relation (Dainotti et al., 2008, 2011, 2015, 2017), a roughly inversely proportional relationship between the rest-frame time at the end of the plateau phase ($T_{a}/(1+z)$) and its corresponding X-ray luminosity ($L_{a}$). Later, Dainotti et al. (2013) showed via the use of the Efron and Petrosian method (Efron & Petrosian (1992)), this relation is intrinsic and not due to selection biases nor the redshift evolution. The Dainotti Relation has also been observed in radio wavelengths (Dainotti et al., 2020; Levine et al., 2022). It has also been extended in three dimensions in X-ray, optical (Dainotti et al., 2022a), and $\gamma$-rays (Dainotti et al., 2021b), where the peak prompt luminosity ($L_{peak}$) has been added to the two-dimensional Dainotti relation (Dainotti et al., 2016, 2017, 2020).
@@
The Dainotti relationships (both in two and three dimensions) were a subject of cosmological investigations, and have been utilized as valuable tools (Cardone et al., 2009, 2010; Dainotti et al., 2013; Postnikov et al., 2014; Cao et al., 2021, 2022b, 2022a; Dainotti et al., 2022b, 2023; Bargiacchi et al., 2023; Dainotti et al., 2023). The Dainotti relations have been shown to provide compatible results for $\Omega_{M}$ (the matter density parameter) under the assumption of flat $\Lambda$CDM model when used in conjunction with supernovae (SNe) Ia (Dainotti et al., 2022b). This approach also has the advantage of expanding the distance ladder up to $z$ = 5, which is much higher than the most distant SNe Ia observed up to date ($z$ = 2.9) (Pierel et al., 2024). Wang et al. (2022) further extended the Dainotti relation up to $z=5.91$. They also standardized the LGRB sample using a method similar to the Phillips Correlation (Phillips, 1993). This standardization helped Wang et al. (2022) achieve a lower scatter on the rescaled luminosity, thus highlighting the importance of the plateau phase as a possible standard candle.
@@
Another notable effort in establishing a fundamental GRB relation is the so-called Combo relation (Izzo et al., 2015), joining $L_{a}$, $T_{a}/(1+z)$, powerlaw decay of afterglow, and intrinsic spectral peak energy of prompt (Ep,i). Izzo et al. (2015) demonstrated, with a limited sample of 5 GRBs, a minimization of the dependency on SNe Ia for calibrating GRBs. They use a two-step calibration process for the Combo relation and measured $\Omega_{M}=0.29^{+0.23}_{-0.15}$ in $\Lambda$CDM.
@@
Previously, there have been several other efforts aiming to enhance the count of GRBs with known redshifts by exploring correlations between distance-independent parameters (such as peak flux and afterglow plateau duration) and distance-dependent GRB characteristics (such as prompt emission and peak luminosity) to estimate pseudo-redshifts for GRBs lacking observed values(Reichart, 2001; Atteia, 2003, 2005; Yonetoku et al., 2004; Dainotti et al., 2011). Only very few GRBs can reach the small uncertainty (5%) (Guiriec et al., 2016). Specifically, when using a relation between the rest-frame $T_{a}$ and its luminosity, only 28% of cases have uncertainties $\Delta z/z_{obs}<0.90$ (Dainotti et al., 2011). If the redshift is extracted by employing the Amati correlation (Amati et al., 2002) between $E_{peak}$ (the peak in the $\nu F_{\nu}$ spectrum) and the energy emitted isotropically during the prompt emission, the Pearson coefficient ($r$) of correlation between $z_{obs}$ and predicted redshift ($z_{pred}$) is $0.67$. Amati (2006) provided $z_{pred}$ for 17 GRBs, and these are accurate within a factor of 2.
Another often overlooked issue is the inherent inaccuracies of these pseudo-redshifts, arising from their non-linear dependence on the luminosity distance ($D_{L}$).
@@
The major problem in the forward-fitting relationships is the dependence on cosmology. Indeed, one of the parameters of this relationship depends on the $D_{L}$, which itself depends on cosmological parameters. We also would like to stress that it is important to have a method that is not dependent on a given cosmological model; otherwise, we incur the so-called ”circularity problem”. Indeed, it would be highly circular to use the pseudo-redshifts derived from the forward fitting method, which assumes a given cosmology to then derive cosmological parameters. Even if the goal of the pseudo-redshift is not to derive the cosmological parameters, we avoid being biased toward a particular or a given cosmological model for the analysis of the luminosity function and density rate evolution here. Although, the errorbars for the luminosity function are not negligible, in the procedure of the selection biases detailed in (Dainotti et al., 2024b) if the cosmology had been inferred, we would have had a different number of associated sets (described in the appendix of that paper), because of the choice of a different limiting luminosity. This, hence, will also affect the density rate evolution.
@@
Similar work has been done by Ukwatta et al. (2016); Aldowma & Razzaque (2024). However, here, we use the plateau features, whose correlations have less intrinsic scatter than prompt ones. This can ensure more stable ML results.

### v3

@@ Abstract
Context. Gamma-ray bursts (GRBs), which have been observed at redshifts as high as 9.4, could serve as valuable probes for investigating the distant Universe. However, using them in this manner necessitates an increase in the number of GRBs with determined redshifts, as currently only 12% of them have known redshifts due to observational biases.
Aims. We aim to address the shortage of GRBs with measured redshifts to enable full realization of their potential as valuable cosmological probes.
Methods. Following our previous approach, in this work we take a further step to overcome this issue by adding 30 more GRBs to our ensemble supervised machine learning training sample, representing an increase of 20%, which will help us obtain more accurate pseudo-redshifts. In addition, we have built a freely accessible and user-friendly web application that infers the redshift of long GRBs (LGRBs) with plateau emission using our machine learning model. The web app is the first of its kind for such a study and will allow the community to obtain pseudo-redshifts by entering the GRB parameters into the app.
Results. Through our machine learning model, we successfully estimated redshifts for 276 LGRBs using X-ray afterglow parameters detected by the Neil Gehrels Swift Observatory and increased the sample of LGRBs with known redshifts by 110%. We also performed Monte Carlo simulations to demonstrate the future applicability of this research.
Conclusions. The results presented in this work will enable the community to increase the sample of GRBs with known pseudo-redshifts. This can help address many outstanding issues, such as GRB formation rate, luminosity function, and the true nature of low-luminosity GRBs, and it can enable the application of GRBs as standard candles.
@@ 1 Introduction
Gamma-ray bursts (GRBs) are the brightest and most powerful explosion events after the Big Bang and are observed across a wide range of redshifts, from 0.0085 (Galama et al., 1998) to 8.2 and 9.4 (Tanvir et al., 2008; Cucchiara et al., 2011). This vast range of redshifts ($z$) makes GRBs ideal probes for studying the early Universe and tracking its evolution to gain insights into the composition, expansion rate, and other key aspects. GRBs, generally detected in $\gamma$-rays, X-rays, and occasionally in optical wavelengths, are classified into two types based on their duration, which is characterized by the $T_{90}$ parameter. The parameter $T_{90}$ is the time span during which the GRB emits 90% of its total observed fluence in $\gamma$-rays. The GRBs with $T_{90}<2s$ are classified as short GRBs (SGRBs), and those with $T_{90}>2s$ are classified as long GRBs (LGRBs). The LGRBs are produced by core collapse supernovae (Mazets et al., 1981; Kouveliotou et al., 1993; Paczyński, 1998; Woosley & Bloom, 2006), while SGRBs are produced by neutron star–neutron star or neutron star–black hole mergers (Narayan et al., 1992; Duncan & Thompson, 1992; Usov, 1992; Thompson, 1994; Nakar, 2007; Goldstein et al., 2017a, b; Abbott et al., 2017).
@@
In addition to these two main categories of GRBs, numerous other sub-classes have been identified in the literature. These classes are based only on observational properties such as the $T_{90}$ duration as well as spectral and light curve (LC) properties, and they include classes such as SGRBs with extended emission (SEE; Norris et al. (2000); Norris & Bonnell (2006); Levan et al. (2007); Norris et al. (2010)), where $T_{90}$ is greater than two seconds but with similar spectral properties to SGRBs, Very Long GRBs (VLGRBs, Levan et al. (2014)) and Ultra Long GRBs (ULGRBs, Stratta et al. (2013); Zhang et al. (2014); Gendre et al. (2019), where $T_{90}$ is greater than $500$ or $1000$ seconds, respectively. Zhang et al. (2007) unified these various classes based on their GRB progenitors. According to their prescription, Type I GRBs comprise classes that are SGRBs, SEEs, and other GRBs similar to SGRBs and occur due to compact binary mergers (Nakar, 2007; Abbott et al., 2017). Type II GRBs comprise LGRBs, VLGRBs, ULGRBs, and other similar classes with core-collapse supernova progenitors (the so-called collapsar events) (Woosley, 1993; Woosley et al., 1993). Irrespective of these efforts, an exhaustive classification of GRBs has not yet been found, and this is an active field of research.
@@
Phenomenologically, the GRB LC from across all the classes can be separated into two distinct phases. The prompt emission is the main emission, and it is characterized by a burst of high-energy photons spanning $\gamma$-rays to hard X-rays that sometimes extends into optical wavelengths (Vestrand et al., 2005; Beskin et al., 2010; Gorbovskoy et al., 2012; Vestrand et al., 2014). Following the prompt emission, the afterglow is a long-lasting, multiwavelength emission powered by the interaction of the GRB ejecta with the surrounding medium (van Paradijs et al., 1997; Costa et al., 1997; Piro et al., 1998). It is detected across various wavelengths, such as X-rays, optical, and occasionally even radio.
@@
A peculiar trait of GRBs, uncovered by the Neil Gehrels Swift observatory (hereafter Swift, Gehrels et al. (2004)), is the presence of a flat region in the LC, the so-called plateau emission, which is characterized by almost constant flux values (Nousek et al., 2006; Zhang et al., 2006; O’Brien et al., 2006; Sakamoto et al., 2007; Liang et al., 2007b; Dainotti et al., 2008; Zaninoni et al., 2013; Rowlinson et al., 2014a). The end of the plateau is considered to be a point in the LC’s afterglow where the luminosity begins to decay with time as a power law. The plateau feature has been found in 42% of X-ray afterglows (Evans et al., 2009; Li et al., 2018). In our recent estimate, the percentage of GRBs with an observable plateaus is 38.5%.
@@
Currently, the primary hurdle in the study of the LGRB population is the lack of a large GRB sample with known redshift. Quick localization and spectral data are required to determine the redshift of GRBs. Notably, Swift allows for this quick detection as well as multi-wavelength follow-up in $\gamma$-rays, X-rays, and ultraviolet/optical. Swift has three telescopes: the Burst Alert Telescope (BAT; Burrows et al. (2005)), the X-ray Telescope (XRT; Barthelmy et al. (2005)), and the Ultraviolet/Optical Telescope (UVOT; Roming et al. (2005)). It uses XRT for precise localization, and with UVOT, it obtains spectra and sometimes redshifts. Swift’s rapid localization abilities have contributed significantly to exploring the high-$z$ Universe.
@@
Although Swift triggers many observation programs, only 26% (431) of its 1656 GRBs have known spectroscopic redshifts (as of 20 August 2024). The Fermi Telescope has observed more than 3700 GRBs with its Large Area Telescope (LAT; Ajello et al. (2019)) and the Fermi Gamma-ray Burst Monitor (GBM; Meegan et al. (2009)). Currently, only 12% of the GRBs detected by Swift and Fermi have spectroscopic redshift ($z_{obs}$), and approximately 90% of them were observed by Swift. Measuring redshift, especially high-$z$, is challenging due to limited telescope time and the difficulty of securing time for follow-up observations using large aperture telescopes. Therefore, determining the $z$ of GRBs via alternative methods is crucial.
@@
Correctly determining the GRB luminosity function (LF) is one of the main goals that can be achieved by increasing the number of GRBs with known redshifts. The LF provides the number of GRBs per unit luminosity and is vital for understanding the properties of the GRB population, such as their energy release and emission mechanisms. Another principal objective is determining the cosmic GRB formation rate (GRBFR). The GRBFR measures the frequency of GRBs across cosmic time.
@@
Petrosian et al. (2015); Petrosian & Dainotti (2024) have found a discrepancy at low redshift ($10<z<$ 1) between the LGRB formation rate and the star formation rate (SFR). This discrepancy, which has been observed by several research groups independently, remains an active area of investigation, and addressing it requires a bigger sample of GRBs with known redshift, thus highlighting the importance of our study.
@@
The correct determination of the GRB rate for both LGRBs and SGRBs is hampered by the broad range of GRB physical features; therefore, a precise distribution of these features and their evolution over cosmic time is necessary. In Dainotti et al. (2024b), we investigated this issue with a sample of 179 GRBs with optical LCs, where we determined the luminosity and formation rate evolutions and the general shape of the LF after correcting for evolution. We observed the average LGRB’s rate at a low-$z$ ($z<1$) to be 10.46 Gpc-3 $yr^{-1}$, which is roughly three times more than Hopkins & Beacom (2006) SFR (3.7 Gpc-3 $yr^{-1}$). This LGRB rate is an actual rate where the normalization is in physical units, which contrasts with the arbitrary normalization used in most of the current literature.
@@
To address the challenges outlined above, we propose a machine learning (ML)-based approach for estimating the redshift of GRBs. This work is a follow-up of Dainotti et al. (2024c) and Dainotti et al. (2024b), and we utilize a sample that is 20% larger than the sample in Dainotti et al. (2024c) to train our ML models. In addition, we extend the generalization sample by 35%. This sample is the set of GRBs without known redshifts, and we predict the redshift of 276 GRBs using our trained ML model. Furthermore, we release a user-friendly web application that provides pseudo-redshifts for new GRBs built on the ML models described in this paper.
@@
The paper is structured as follows. In Sec. 2, we discuss the advantages of using an ML approach compared to using linear relations for the redshift estimation. In Sec. 3, we describe our data set. In Sec. 4, we describe the functionalities and various modules of our web app as well as details of the ML models being used. We also describe the Monte Carlo (MC) simulation performed to understand the viability of our ML models in relation to future data. Sec. 5 shows the results obtained from our analysis and the usability of the web app. Here, we also present the future applicability of our ML approach. In Sec. 6, we discuss the relevance of our results. Finally, we summarize and conclude in Sec. 7.
2 Advantage of machine learning versus linear relationships
@@
Expanding the GRB sample with redshift provides a significant benefit, notably the possibility of using them as standardized candles, objects with known intrinsic luminosities, or with luminosities that can be derived via established relationships among GRB properties. The adoption of empirical relations between the distance-dependent and intrinsic properties of GRBs enables cosmological studies in unexplored redshift ranges. One of the earliest efforts is the Dainotti Relation (Dainotti et al., 2008, 2011, 2015, 2017), a roughly inversely proportional relationship between the rest-frame time at the end of the plateau phase ($T_{a}/(1+z)$) and its corresponding X-ray luminosity ($L_{a}$). Later, Dainotti et al. (2013) showed via the use of the Efron and Petrosian method (Efron & Petrosian (1992)), this relation is intrinsic and not due to selection biases nor the redshift evolution. The Dainotti Relation has also been observed in radio wavelengths (Dainotti et al., 2020; Levine et al., 2022). It has also been extended in three dimensions in X-ray, optical (Dainotti et al., 2022), and $\gamma$-rays (Dainotti et al., 2021b), where the peak prompt luminosity ($L_{peak}$) has been added to the two-dimensional Dainotti relation (Dainotti et al., 2016, 2017, 2020).
@@
The Dainotti relationships (both in two and three dimensions) were a subject of cosmological investigations, and have been utilized as valuable tools (Cardone et al., 2009, 2010; Dainotti et al., 2013; Postnikov et al., 2014; Cao et al., 2021, 2022b, 2022a; Dainotti et al., 2023a, b; Bargiacchi et al., 2023; Dainotti et al., 2023). The Dainotti relations have been shown to provide compatible results for $\Omega_{M}$ (the matter density parameter) under the assumption of flat $\Lambda$CDM model when used in conjunction with supernovae (SNe) Ia (Dainotti et al., 2023a). This approach also has the advantage of expanding the distance ladder up to $z$ = 5, which is much higher than the most distant SNe Ia observed up to date ($z$ = 2.9) (Pierel et al., 2024). Wang et al. (2022) further extended the Dainotti relation up to $z=5.91$. They also standardized the LGRB sample using a method similar to the Phillips Correlation (Phillips, 1993). This standardization helped Wang et al. (2022) achieve a lower scatter on the rescaled luminosity, thus highlighting the importance of the plateau phase as a possible standard candle.
@@
Another notable effort in establishing a fundamental GRB relation is the so-called Combo relation (Izzo et al., 2015), joining $L_{a}$, $T_{a}/(1+z)$, power-law decay of afterglow, and intrinsic spectral peak energy of prompt (Ep,i). Izzo et al. (2015) demonstrated, with a limited sample of 5 GRBs, a minimization of the dependency on SNe Ia for calibrating GRBs. They use a two-step calibration process for the Combo relation and measured $\Omega_{M}=0.29^{+0.23}_{-0.15}$ in $\Lambda$CDM.
@@
Previously, there have been several other efforts aiming to enhance the count of GRBs with known redshifts by exploring correlations between distance-independent parameters (such as peak flux and afterglow plateau duration) and distance-dependent GRB characteristics (such as prompt emission and peak luminosity) to estimate pseudo-redshifts for GRBs lacking observed values(Reichart, 2001; Atteia, 2003, 2005; Yonetoku et al., 2004; Dainotti et al., 2011). Very few GRBs can reach the small uncertainty (5%) (Guiriec et al., 2016). Specifically, when using a relation between the rest-frame $T_{a}$ and its luminosity, only 28% of cases have uncertainties $\Delta z/z_{obs}<0.90$ (Dainotti et al., 2011). If the redshift is extracted by employing the Amati correlation (Amati et al., 2002) between $E_{peak}$ (the peak in the $\nu F_{\nu}$ spectrum) and the energy emitted isotropically during the prompt emission, the Pearson coefficient ($r$) of correlation between $z_{obs}$ and predicted redshift ($z_{pred}$) is $0.67$. Amati (2006) provided $z_{pred}$ for 17 GRBs, and these are accurate within a factor of 2.
Another often overlooked issue is the inherent inaccuracies of these pseudo-redshifts, which arise from their non-linear dependence on the luminosity distance ($D_{L}$).
@@
The major problem in the forward-fitting relationships is the dependence on cosmology. Indeed, one of the parameters of this relationship depends on the $D_{L}$, which itself depends on cosmological parameters. We also would like to stress that it is important to have a method that is not dependent on a given cosmological model; otherwise, we incur the so-called ”circularity problem”. Indeed, it would be highly circular to use the pseudo-redshifts derived from the forward fitting method, which assumes a given cosmology to then derive cosmological parameters. Even if the goal of the pseudo-redshift is not to derive the cosmological parameters, we avoid being biased toward a particular or a given cosmological model for the analysis of the LF and density rate evolution here. Although, the error bars for the LF are not negligible, in the procedure of the selection biases detailed in (Dainotti et al., 2024b) if the cosmology had been inferred, we would have had a different number of associated sets (described in the appendix of that paper), because of the choice of a different limiting luminosity. Hence, this will also affect the density rate evolution.
@@
Similar work has been done by Ukwatta et al. (2016); Aldowma & Razzaque (2024). However, here we use the plateau features, whose correlations have less intrinsic scatter than prompt ones. This can ensure more stable ML results.